find d2y/dx2 where y=log(x^2/e^2)

Grade:12

2 Answers

Rinkoo Gupta
9 years ago
y=log(x2/e2)
diff. w.r. to x , we get
dy/dx=1/(x2/e2) . (2x/e2)
=2/x

diff. w.r. to x again
d2y/dx2 =-2/x2
Thanks & Regards
Rinkoo Gupta
AskIITians Faculty

Rinkoo Gupta
9 years ago

y=log(x2/e2)

diff. w.r. to x , we get

dy/dx=1/(x2/e2) . (2x/e2)

=2/x

diff. w.r. to x again

d2y/dx2=-2/x2

Thanks & Regards

Rinkoo Gupta

AskIITians Faculty